OpenStudy (anonymous):

HELP ME PLEASE!!!! The volume of a box is 1344 cubic inches. What is the volume of the box in cubic feet?

OpenStudy (aripotta):

just divide 1344 by 12 because there's 12 inches in a foot

OpenStudy (anonymous):

oh okayy thanks.

OpenStudy (anonymous):

1 inch = 0.0833333 foot 1 inch^3 = 0.083333^3 foot^3 = 0.0005787 foot^3 1344 * 0.0005787 = 0.7778 foot^3

OpenStudy (anonymous):

thank you too.